OpenStudy (anonymous):

Prove (1-sin(x))(sec(x)+tan(x))=cos(x)

OpenStudy (anonymous):

(1-sin x)(sec x+tan x)=(1-sin x)(1/cos x+sin x/cos x)=(1-sin x)(1+sin x)/cos x=(1-sin ^2 x)/cos x=cos ^ 2x/cos x=cos x .@renap

OpenStudy (anonymous):

@renap